# A Wavefield Correlation Approach to Improve Sound Speed Estimation in Ultrasound Autofocusing

### v1

@@ V-A Sound Speed Estimation and Aberration Correction Enhancements
A major difference in the sound speed estimates from the WFC model is the accuracy in the sound speed values and structure compared to DAS with a straight-ray travel path assumption. By simulating the wave propagation through a heterogeneous medium, the refraction and diffraction of the waves is included in the wave propagation model, which is more accurate than a straight-ray assumption that does not include refraction or diffraction. This effectively results in more accurate beamforming in heterogeneous media, especially with larger sound speed variations, and more accurate estimation of the sound speed structure, as exemplified by the shape of the inclusions in Figure 4. The ground truth sound speeds were not measured for this inclusion phantom, but the material composition formulas used in Ali et al. [4] suggest the target sound speeds were 1650 and 1515 m/s for the inclusions and background, respectively. Although these sound speeds cannot be guaranteed, a higher sound speed estimate can be observed in both the inclusion and background of this phantom when using WFC compared to straight-ray DAS. While sound speed estimates from the straight-ray model can result in a visible improvement to the final aberration-corrected image, a more accurate sound speed map can further improve the image, as well as provide more reliable sound speed estimates for use in quantitative biomarkers of disease. Additionally, due to the spatiotemporal matched filtering from the correlation of an idealized forward-propagated transmit wavefield and the backward-propagated actual received wavefield, the effect of diffuse clutter can be reduced using WFC, which can improve the calculation of image quality metrics which would otherwise be adversely affected by reverberation noise. This results in more accurate sound speed estimations in cluttered scenarios, such as the reverberation shown in the lumen of the vasculature in Fig. 5. However, this spatiotemporal matched filtering only occurs with the one-way propagation simulation method used here. A two-way propagation model, such as k-Wave or Fullwave, will match the reverberation signal from propagation of the reflected waves and will thus not cancel multiple reflections.
@@
Some small artifacts worth noting in some of the WFC sound speed estimation results are the corner artifacts in the two simulation cases in Figure 2 and the sound speed dips on the edges of the meat-layer phantom in Figure 3. These features do not consistently appear for all sources of data, but they could potentially be mitigated through different regularization parameters or setup, changes to the field of view covered by the points in the loss calculation grid, or different initialization parameters.
V-B Optimization Approach
An advantage of auto-differentiation to optimize sound speed through gradient descent is the relative ease through which arbitrary loss criteria without an explicitly calculated gradient can be backpropagated to update a variable of interest, as long as all operations taken to calculate the loss with respect to that variable are backwards differentiable in software. Wavefield correlation as described in this work is a fully differentiable approach that accounts more accurately for wave propagation in a heterogeneous medium. Calculating Eikonal time delays is an alternative approach that accounts for refraction compared to the simpler straight ray approximation, but typical Eikonal equation solvers are not differentiable [31]. Additionally, the CMPE metric, which can be calculated in a differentiable way using separate sub-aperture split-step angular spectrum simulations, results in better sound speed estimates compared to other image focus metrics, as shown in [32] and [13]. However, this optimization approach can have several disadvantages, some of which are similar to other sound speed estimators for pulse-echo ultrasound. Tracking all operations and necessary variable values to perform software auto-differentiation requires a considerable amount of memory, and can easily require high-end GPUs with more vRAM when utilizing GPU code acceleration. While DAS beamformers can benefit more from the GPU acceleration and have fewer operations to track, the angular spectrum simulation used for WFC must run sequentially in depth, resulting in longer runtimes, and more computational steps result in higher memory needs. However, this style of optimization at this stage of development must be performed offline anyway due to the gradient-descent-style iterative loss minimization. To reduce the runtime by requiring fewer iterative steps, a heterogeneous local sound speed map, obtained from a real-time sound speed estimator, could be used to initialize the optimization closer to the actual sound speed. Additionally, loss metrics that are faster to calculate than CMPE could be used for some or all parts of the iteration process.
Similar to other sound speed estimators, a few components in the algorithm require manual tuning or adjustment. Regularization has a large impact on the morphology of the final result due to the ill-posedness of the pulse-echo local estimation problem. Simply optimizing for an image metric by itself often does not result in a physically possible sound speed map, especially in cluttered data, due to the large, rapid variations or extreme values that will develop in the sound speed map. Regularization, therefore, must be added to the optimized loss, with the exact weight tuned for different media to enforce a balance in realistic smoothness and sound speed variation. Additionally, the stopping criterion is not standardized in this work and was set to a number of iterations for which the sound speed estimation settled to stable values. To more adaptively stop the iteration process, a stopping criterion can be defined based on the CMPE, which could further reduce the runtime.
V-C Initialization and Pulse-Echo Depth-Velocity Ambiguity
@@
Due to the extreme non-convexity and ill-posedness of the optimization problem, the initialization greatly affects the global (mean) sound speed and, consequently, the bias in the final sound speed map. The effect of the initialization on the final global sound speed offset has been demonstrated for other pulse-echo local sound speed estimators, like Figure 5 in [4]. While the exact global sound speed, and thus the spatial locations of structures in the beamformed image, heavily depends on the constant sound speed used for initialization, aberration is still corrected to a similar degree in the final image using the estimated sound speed. Different initializations, coupled with differences in the field of view used, can also affect the final sound speed estimation results, as seen when comparing the fabricated inclusion phantoms in Figure 4 and the same phantom estimates in [32] with a different field of view and initialization parameters. This phenomenon is caused by depth-velocity ambiguity in pulse-echo measurements, where the exact location of objects in a medium cannot be determined if the sound speed is unknown. This ambiguity is presented in theoretical detail in [24], and similar occurrences of depth-velocity ambiguity in iterative sound speed estimation heavily influenced by initialization are discussed in [4] and [5]. Although depth-velocity ambiguity can make the exact sound speed values estimated less reliable to directly use for downstream applications, the relative change in sound speed in the map is still useful for applications like aberration correction. Regardless of the location of features due to the global sound speed offset, resolution and contrast still improve from accounting for the relative time delay changes caused by changing sound speed in media.
@@
An approach to reduce the effects of initialization is calibration (which can additionally account for modeling errors). Tomographic inversion approaches to estimating sound speed use phase shift calibration for different starting sound speeds to account for modeling errors [16]. For an iterative sound speed estimation approach, [32] instead utilizes an additional scalar time offset that is calibrated for individual sources of data. This time offset essentially accounts for portions of the wave propagation process that are not modeled, such as the transducer lens and 3D propagation. This time offset is chosen to minimize the sound speed bias for known homogeneous sound speed data for each data collection scenario, requiring separate calibration for both simulation and different scanned media to reduce estimation bias. While calibration is more reliable in reducing the global bias, it either requires pre-calibration for each imaging scenario and starting sound speed, or knowing the sound speed in a medium, which could complicate standard clinical in vivo imaging applications. In this work, both the initial sound speed and a time offset are both swept prior to the initial beamforming iteration to find the pair of values that optimizes the CMPE. While this can result in lower estimation bias than simply sweeping for a starting sound speed, the initialization can have greater variation based on factors such as the exact medium imaged, field of view, or number of elements used during beamforming.
@@
However, to better address the global sound speed bias without calibration, future approaches would either need to overcome the depth-velocity ambiguity, or reliably initialize the sound speed and accurately model the exact wave propagation for scanner data. Better modeling of wave propagation through the transducer lens, as in [36], could reduce the need for applying a time offset to account for propagation inaccuracies. Additionally, initialization could be posed as a separate problem to optimize by itself due to the strong influence of the initial sound speed on the final global sound speed of the estimate. For example, a hyperparameter tuning approach could be utilized to find the optimal parameters for initialization, as well as the rest of the optimization process.
@@ VI Conclusion
We present an approach to local sound speed estimation that improves estimates in challenging heterogeneous media, especially cases with reverberation clutter and large changes in sound speed. The proposed estimator optimizes sound speed by implementing the WFC beamformer, which involves simulating transmit and receive wavefield propagation, using auto-differentiation software. A total-variation regularized CMPE loss is calculated from beamformed data, and a local sound speed map is iteratively updated using gradient descent, with the backpropagated gradient automatically computed through software. We compare sound speed estimates and aberration correction for this beamformer with a previously-published straight-ray DAS approach, showing larger estimated sound speed variation between different media regions, decreased estimation error, and improved image resolution and contrast. These results show promise for extending pulse-echo sound speed estimation to more challenging in vivo imaging situations with significant clutter and sound speed changes.
@@ Acknowledgment
The authors would like to thank the Dahl Lab beamforming group for their helpful discussions and insight.

### v2

@@ V-A Sound Speed Estimation and Aberration Correction Enhancements
The major improvements provided by the WFC model with respect to DAS beamforming with a straight-ray model include improved sound speed accuracy and reconstruction of complex tissue structures. By simulating the wave propagation through a heterogeneous medium, refraction and diffraction of the waves are included in the wave propagation model, improving upon our prior DAS straight-ray model. In our prior model [35], we noted that the straight-ray model does not account for refraction or diffraction, resulting in errors in sound speed estimation. However, in the prior model yielded good aberration correction and image quality despite these errors due to the ill-posedness of the problem; that is, there are many solutions to the sound speed estimates that resulted in low CMPE. In our present WFC model, the use of improved modeling of wave propagation reduces the attribution of model error, instead of aberration, to the CMPE loss funciton. This is what leads to more accurate estimation of the speed of sound, especially when larger sound speed variations are present, as demonstrated by the shape of the inclusions in Figures 2(a) and 4. While the ground truth sound speeds were not availabe for this inclusion phantom, the material composition formulas used in Ali et al. [4] suggest the target sound speeds were 1650 and 1515 m/s for the inclusions and background, respectively. Although these sound speeds cannot be guaranteed [^1] Anecdotally, it has been observed that gelatin-alcohol phantoms of different composition will diffuse into each other, resulting in a change in sound speed., a higher sound speed estimate can be observed in both the inclusion and background of this phantom in the WFC model compared to DAS straight-ray model.
@@
While sound speed estimates from the straight-ray model can result in visible improvement to the final aberration-corrected images, a more accurate sound speed map can further improve image quality, as well as provide more reliable sound speed estimates for use in quantitative biomarkers of disease. Additionally, due to the spatiotemporal matched filtering from the correlation of an idealized forward-propagated transmit wavefield and the backward-propagated actual received wavefield, the effect of diffuse clutter can be reduced using WFC, which can improve the fidelity of the ultrasound signals that would otherwise be adversely affected by reverberation noise. This results in more accurate sound speed estimations and imporved image quality in cluttered scenarios, such as the reverberation shown in the lumen of the vasculature in Fig. 5a. However, this spatiotemporal matched filtering only occurs with the one-way propagation simulation method used here. A two-way propagation model, such as k-Wave or Fullwave, will match the reverberation signal from propagation of the reflected waves and is not expected to cancel the multiple reflections.
@@
There are a few anomalies worth noting with the sound speed estimates in the WFC model. First, the small artifacts at the corners of the estimates in the two simulation cases in Figure 2 and the sound speed dips on the edges of the meat-layer phantom in Figure 3 do not consistently appear for all sources of data, but are also not uncommon. These artifacts can potentially be mitigated through different regularization parameters, changes to the field of view in the loss calculation or speed of sound grids, or different initialization parameters. Second, the worse sound speed estimates found in the bottom layer of Figure 2(b) compared to the DAS model was an uncommon result. In the majority of simulation and experimental data containing heterogeneous distributions of speed of sound, we found the WFC model to more accurately estimate the speed of sound in terms of bias and MAE.
V-B Initialization and Pulse-Echo Depth-Velocity Ambiguity
@@
Due to the extreme non-convexity and ill-posedness of the optimization problem, initialization greatly affects the global (mean) sound speed and, consequently, the bias in the final sound speed map. The effect of the initialization on the final global sound speed offset has been demonstrated for other pulse-echo local sound speed estimators, such as Figure 10 in [17] and Figure 5 in [4]. While the exact global sound speed—and thus the spatial locations of structures in the beamformed image—heavily depends on the constant sound speed used for initialization, aberration is still corrected to a similar degree in the final image using the estimated sound speed. Different initializations, coupled with differences in the field of view, can also affect the final sound speed estimation results, as seen when comparing the fabricated inclusion phantoms in Figure 4 and the same phantom estimates in [35] with a different field of view and initialization parameters. This phenomenon is caused by depth-velocity ambiguity in pulse-echo measurements, where the exact location of objects in a medium cannot be determined if the sound speed is unknown. This ambiguity is presented in theoretical detail in [27], and similar occurrences of depth-velocity ambiguity in iterative sound speed estimation heavily influenced by initialization are discussed in [4] and [5]. Although depth-velocity ambiguity can make the estimated sound speed values less reliable for direct use in downstream applications, the relative change in sound speed in the map is still useful for applications like aberration correction. Regardless of the location of features due to the global sound speed offset, resolution and contrast still improve from accounting for the relative time delay changes caused by changing sound speed in media.
@@
An approach to reduce the effects of initialization is calibration (which can additionally account for modeling errors). Tomographic inversion approaches to estimating sound speed use phase shift calibration for different starting sound speeds to account for modeling errors [17]. For an iterative sound speed estimation approach, [35] instead utilizes an additional scalar time offset that can be calibrated for individual sources of data. This time offset essentially accounts for portions of the wave propagation process that are not modeled, such as the transducer lens and 3D effects. This time offset is chosen to minimize the sound speed bias for known homogeneous sound speed data for each data collection scenario, requiring separate calibration for both simulation and different scanned media to reduce estimation bias. While calibration approaches are more reliable in reducing the global bias, it either requires pre-calibration for each imaging scenario and starting sound speed, or knowing the sound speed in a medium, which could complicate standard clinical in vivo imaging applications. In this work, both the initial sound speed and a time offset are swept prior to the initial beamforming iteration to find the pair of values that optimizes the CMPE. While this can result in lower estimation bias than simply sweeping for a starting sound speed, the initialization can have greater variation based on factors such as the exact medium imaged, field of view, or number of elements used during beamforming.
@@
However, to better address the global sound speed bias without calibration, future approaches would either need to overcome the depth-velocity ambiguity, or reliably initialize the sound speed and more accurately model the wave propagation in ultrasound scanner data. Better modeling of wave propagation through the transducer lens, as in [39], could reduce the need for applying a time offset to account for propagation inaccuracies. Additionally, initialization could be posed as a separate problem to optimize by itself due to the strong influence of the initial sound speed on the final global sound speed of the estimate. For example, a hyperparameter tuning approach could be utilized to find the optimal parameters for initialization, as well as the rest of the optimization process.
V-C Optimization Approach
An advantage of auto-differentiation to optimize sound speed through gradient descent is the relative ease through which arbitrary loss criteria without an explicitly calculated gradient can be backpropagated to update a variable of interest, as long as all operations taken to calculate the loss with respect to that variable are backwards differentiable. Wavefield correlation as described in this work is a fully differentiable approach that accounts more accurately for wave propagation in a heterogeneous medium. Calculating Eikonal time delays is an alternative approach that accounts for refraction compared to the simpler straight ray approximation, but typical Eikonal equation solvers are not differentiable [34]. At this time, we have found that the CMPE metric, which can be calculated in a differentiable way using separate sub-aperture split-step angular spectrum simulations, results in consistently better sound speed estimates compared to other image focusing metrics we have explored, such as those shown in [35] and [14].
The proposed optimization approach can have several disadvantages, some of which are similar to other sound speed estimators for pulse-echo ultrasound. Tracking all operations and necessary variable values to perform software auto-differentiation requires a considerable amount of memory, and can easily require high-end GPUs with more vRAM when utilizing GPU code acceleration. While DAS beamformers can benefit more from the GPU acceleration and have fewer operations to track, the angular spectrum simulation used for WFC must run sequentially in depth, resulting in longer runtimes, and more computational steps result in higher memory needs. However, the style of optimization at this stage of development must be performed offline anyway due to the gradient-descent-style iterative loss minimization. To reduce the runtime by requiring fewer iterative steps, a heterogeneous local sound speed map obtained from a real-time sound speed estimator could be used to initialize the optimization closer to the actual sound speed. Additionally, loss metrics that are faster to calculate than CMPE could be used for some or all parts of the iteration process, although this may come at the expense of performance.
Similar to other sound speed estimators, a few components in the algorithm require manual tuning or adjustment. Regularization has a large impact on the morphology of the final result due to the ill-posedness of the pulse-echo local estimation problem. Simply optimizing for an image metric by itself often does not result in a physically plausible sound speed map, especially in cluttered data, due to the large rapid variations or extreme values that will develop in the sound speed map. Regularization, therefore, must be added to the optimized loss with the exact weight tuned for different media to enforce a balance in realistic smoothness and sound speed variation. Additionally, the stopping criterion is not standardized in this work and was set to a fixed number of iterations for which the sound speed estimation settled to stable values. Instead, an adaptive stopping criterion can be devised in the future based on the loss function, which could further reduce the runtime.
@@ VI Conclusion
We present a method for local sound speed imaging that improves estimation in challenging heterogeneous media compared to a state-of-the-art approach, especially in cases with reverberation clutter and large changes in sound speed. The proposed estimator optimizes sound speed by implementing the WFC beamformer—which involves simulating transmit and receive wavefield propagation—using auto-differentiation software. A total-variation regularized CMPE loss is calculated from beamformed data, and a local sound speed map is iteratively updated using gradient descent, with the backpropagated gradient automatically computed through software. We compare sound speed estimates and aberration correction for the proposed beamformer with a previously-published straight-ray DAS approach, showing larger estimated sound speed variation between different media regions, decreased estimation error, and improved image resolution and contrast. These results are promising for extending pulse-echo sound speed estimation to more challenging in vivo imaging situations with significant clutter and sound speed changes.
@@ Acknowledgment
The authors would like to thank Michael Jaeger, Di Xiao, Pat de la Torre, and Alfred Yu for their helpful discussions and insights regarding the behavior of speed of sound in custom made phantoms.